# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ Appendix B Full Results for Models
| DE | Gemma2 2B IT | 97.221.4 | 92.961.39 | 95.020.48 | 69.76.78 | 86.634.79 | 76.843.12 | 85.931.75 | ||||||||
| Gemma2 9B IT | 96.481.0 | 91.90.9 | 94.130.58 | 65.154.29 | 82.113.76 | 72.532.91 | 83.331.72 | |||||||||
| Qwen2 7B IT | 95.192.1 | 90.180.34 | 92.60.96 | 57.582.14 | 75.467.33 | 65.052.19 | 78.831.56 | |||||||||
| SaigaLlama3 8B | 98.71.19 | 91.590.79 | 95.010.41 | 62.884.08 | 92.885.97 | 74.742.12 | 84.871.22 | |||||||||
| Vikhr 7B IT 0.4 | 94.072.77 | 90.261.15 | 92.11.11 | 58.336.11 | 72.8212.43 | 63.923.46 | 78.012.17 | |||||||||
| Vikhr 7B IT 5.4 | 95.742.52 | 90.82.36 | 93.151.15 | 59.8512.14 | 79.079.93 | 66.918.35 | 80.034.65 | |||||||||
| VikhrGemma 2B IT | 96.481.62 | 94.740.88 | 95.590.66 | 78.034.08 | 84.965.72 | 81.132.42 | 88.361.52 | |||||||||
| DSM | Gemma2 2B IT | 58.649.18 | 63.294.98 | 60.626.33 | 49.079.31 | 44.425.41 | 46.246.49 | 53.435.33 | ||||||||
| Gemma2 9B IT | 56.796.65 | 64.23.12 | 60.164.97 | 52.784.24 | 45.124.14 | 48.543.57 | 54.353.9 | |||||||||
| Qwen2 7B IT | 72.228.21 | 68.876.39 | 70.46.75 | 50.9310.84 | 55.3311.08 | 52.8110.26 | 61.618.32 | |||||||||
| SaigaLlama3 8B | 56.1715.48 | 57.467.16 | 56.1310.59 | 38.8913.98 | 37.439.05 | 37.2110.11 | 46.677.7 | |||||||||
| Vikhr 7B IT 0.4 | 70.9911.4 | 64.122.52 | 67.065.9 | 40.747.64 | 50.7511.53 | 44.074.72 | 55.574.06 | |||||||||
| Vikhr 7B IT 5.4 | 54.328.73 | 55.55.73 | 54.796.93 | 35.197.64 | 34.37.65 | 34.587.06 | 44.696.59 | |||||||||
| VikhrGemma 2B IT | 58.6425.16 | 61.416.24 | 57.6718.04 | 49.0716.79 | 48.6612.71 | 45.686.56 | 51.688.61 | |||||||||
| AL | Gemma2 2B IT | 60.6113.55 | 60.184.51 | 59.668.17 | 53.5110.71 | 54.877.3 | 53.456.06 | 56.565.4 | ||||||||
| Gemma2 9B IT | 46.214.08 | 47.525.48 | 46.794.42 | 40.358.95 | 39.06.04 | 39.67.34 | 43.25.65 | |||||||||
| Qwen2 7B IT | 50.013.38 | 53.2510.65 | 51.411.9 | 50.011.67 | 46.7411.99 | 48.1611.5 | 49.7811.33 | |||||||||
| SaigaLlama3 8B | 53.036.25 | 56.453.35 | 54.323.1 | 51.7511.13 | 48.334.02 | 49.677.13 | 51.993.55 | |||||||||
| Vikhr 7B IT 0.4 | 50.06.94 | 54.875.68 | 52.044.93 | 51.7510.27 | 47.014.37 | 49.016.6 | 50.534.71 | |||||||||
| Vikhr 7B IT 5.4 | 56.068.16 | 57.333.3 | 56.474.96 | 51.756.39 | 50.764.71 | 50.993.89 | 53.733.56 | |||||||||
| VikhrGemma 2B IT | 56.828.6 | 59.32.53 | 57.594.7 | 54.399.92 | 52.13.11 | 52.815.51 | 55.22.95 | |||||||||
| AD | Gemma2 2B IT | 55.8310.17 | 59.215.93 | 56.626.27 | 55.5615.38 | 52.585.69 | 53.369.71 | 54.995.86 | ||||||||
| Gemma2 9B IT | 46.675.53 | 52.034.74 | 49.054.48 | 51.858.28 | 46.54.47 | 48.915.93 | 48.984.5 | |||||||||
| Qwen2 7B IT | 49.1712.72 | 48.12.19 | 47.917.44 | 41.6713.13 | 41.984.13 | 41.087.91 | 44.52.76 | |||||||||
| SaigaLlama3 8B | 53.334.71 | 57.846.15 | 55.223.94 | 55.5612.42 | 51.135.98 | 53.018.89 | 54.125.88 | |||||||||
| Vikhr 7B IT 0.4 | 51.6712.13 | 59.814.42 | 54.667.59 | 61.1111.56 | 53.474.23 | 56.436.19 | 55.554.55 | |||||||||
| Vikhr 7B IT 5.4 | 55.012.58 | 51.918.89 | 53.3210.58 | 44.448.49 | 47.6610.86 | 45.889.46 | 49.69.75 | |||||||||
| VikhrGemma 2B IT | 39.176.72 | 45.934.26 | 42.135.5 | 49.075.93 | 42.093.57 | 45.224.01 | 43.683.91 | |||||||||
| AC | Gemma2 2B IT | 63.74.19 | 58.733.32 | 61.083.43 | 46.935.35 | 52.174.36 | 49.354.71 | 55.223.84 | ||||||||
| Gemma2 9B IT | 61.485.97 | 56.841.85 | 58.852.17 | 44.38.78 | 49.072.0 | 46.215.61 | 52.532.19 | |||||||||
| Qwen2 7B IT | 57.045.97 | 57.363.77 | 57.154.76 | 50.04.02 | 49.734.74 | 49.814.08 | 53.484.25 | |||||||||
| SaigaLlama3 8B | 65.565.7 | 59.023.26 | 62.033.81 | 46.055.84 | 53.164.52 | 49.194.61 | 55.613.74 | |||||||||
| Vikhr 7B IT 0.4 | 60.378.36 | 59.623.49 | 59.785.6 | 51.756.56 | 52.724.82 | 52.04.61 | 55.894.23 | |||||||||
| Vikhr 7B IT 5.4 | 62.968.18 | 61.872.07 | 62.134.3 | 53.957.25 | 55.53.03 | 54.333.58 | 58.232.38 | |||||||||
| VikhrGemma 2B IT | 64.816.21 | 57.772.9 | 60.983.77 | 43.866.39 | 51.44.47 | 47.114.91 | 54.053.57 |

### v3

@@ Appendix B Full Results for Models
| DE | Gemma2 2B IT | 92.96 1.39 | 97.22 1.40 | 95.02 0.48 | 86.63 4.79 | 69.70 6.78 | 76.84 3.12 | 85.93 1.75 | ||||||||
| Gemma2 9B IT | 91.90 0.90 | 96.48 1.00 | 94.13 0.58 | 82.11 3.76 | 65.15 4.29 | 72.53 2.91 | 83.33 1.72 | |||||||||
| Qwen2 7B IT | 90.18 0.34 | 95.19 2.10 | 92.60 0.96 | 75.46 7.33 | 57.58 2.14 | 65.05 2.19 | 78.83 1.56 | |||||||||
| SaigaLlama3 8B | 91.59 0.79 | 98.70 1.19 | 95.01 0.41 | 92.88 5.97 | 62.88 4.08 | 74.74 2.12 | 84.87 1.22 | |||||||||
| Vikhr 7B IT 0.4 | 90.26 1.15 | 94.07 2.77 | 92.10 1.11 | 72.82 12.43 | 58.33 6.11 | 63.92 3.46 | 78.01 2.17 | |||||||||
| Vikhr 7B IT 5.4 | 90.80 2.36 | 95.74 2.52 | 93.15 1.15 | 79.07 9.93 | 59.85 12.14 | 66.91 8.35 | 80.03 4.65 | |||||||||
| VikhrGemma 2B IT | 94.74 0.88 | 96.48 1.62 | 95.59 0.66 | 84.96 5.72 | 78.03 4.08 | 81.13 2.42 | 88.36 1.52 | |||||||||
| DSM | Gemma2 2B IT | 63.29 4.98 | 58.64 9.18 | 60.62 6.33 | 44.42 5.41 | 49.07 9.31 | 46.24 6.49 | 53.43 5.33 | ||||||||
| Gemma2 9B IT | 64.20 3.12 | 56.79 6.65 | 60.16 4.97 | 45.12 4.14 | 52.78 4.24 | 48.54 3.57 | 54.35 3.90 | |||||||||
| Qwen2 7B IT | 68.87 6.39 | 72.22 8.21 | 70.40 6.75 | 55.33 11.08 | 50.93 10.84 | 52.81 10.26 | 61.61 8.32 | |||||||||
| SaigaLlama3 8B | 57.46 7.16 | 56.17 15.48 | 56.13 10.59 | 37.43 9.05 | 38.89 13.98 | 37.21 10.11 | 46.67 7.70 | |||||||||
| Vikhr 7B IT 0.4 | 64.12 2.52 | 70.99 11.40 | 67.06 5.90 | 50.75 11.53 | 40.74 7.64 | 44.07 4.72 | 55.57 4.06 | |||||||||
| Vikhr 7B IT 5.4 | 55.50 5.73 | 54.32 8.73 | 54.79 6.93 | 34.30 7.65 | 35.19 7.64 | 34.58 7.06 | 44.69 6.59 | |||||||||
| VikhrGemma 2B IT | 61.41 6.24 | 58.64 25.16 | 57.67 18.04 | 48.66 12.71 | 49.07 16.79 | 45.68 6.56 | 51.68 8.61 | |||||||||
| AL | Gemma2 2B IT | 60.18 4.51 | 60.61 13.55 | 59.66 8.17 | 54.87 7.30 | 53.51 10.71 | 53.45 6.06 | 56.56 5.40 | ||||||||
| Gemma2 9B IT | 47.52 5.48 | 46.21 4.08 | 46.79 4.42 | 39.00 6.04 | 40.35 8.95 | 39.60 7.34 | 43.20 5.65 | |||||||||
| Qwen2 7B IT | 53.25 10.65 | 50.00 13.38 | 51.40 11.90 | 46.74 11.99 | 50.00 11.67 | 48.16 11.50 | 49.78 11.33 | |||||||||
| SaigaLlama3 8B | 56.45 3.35 | 53.03 6.25 | 54.32 3.10 | 48.33 4.02 | 51.75 11.13 | 49.67 7.13 | 51.99 3.55 | |||||||||
| Vikhr 7B IT 0.4 | 54.87 5.68 | 50.00 6.94 | 52.04 4.93 | 47.01 4.37 | 51.75 10.27 | 49.01 6.60 | 50.53 4.71 | |||||||||
| Vikhr 7B IT 5.4 | 57.33 3.30 | 56.06 8.16 | 56.47 4.96 | 50.76 4.71 | 51.75 6.39 | 50.99 3.89 | 53.73 3.56 | |||||||||
| VikhrGemma 2B IT | 59.30 2.53 | 56.82 8.60 | 57.59 4.70 | 52.10 3.11 | 54.39 9.92 | 52.81 5.51 | 55.20 2.95 | |||||||||
| AD | Gemma2 2B IT | 59.21 5.93 | 55.83 10.17 | 56.62 6.27 | 52.58 5.69 | 55.56 15.38 | 53.36 9.71 | 54.99 5.86 | ||||||||
| Gemma2 9B IT | 52.03 4.74 | 46.67 5.53 | 49.05 4.48 | 46.50 4.47 | 51.85 8.28 | 48.91 5.93 | 48.98 4.50 | |||||||||
| Qwen2 7B IT | 48.10 2.19 | 49.17 12.72 | 47.91 7.44 | 41.98 4.13 | 41.67 13.13 | 41.08 7.91 | 44.50 2.76 | |||||||||
| SaigaLlama3 8B | 57.84 6.15 | 53.33 4.71 | 55.22 3.94 | 51.13 5.98 | 55.56 12.42 | 53.01 8.89 | 54.12 5.88 | |||||||||
| Vikhr 7B IT 0.4 | 59.81 4.42 | 51.67 12.13 | 54.66 7.59 | 53.47 4.23 | 61.11 11.56 | 56.43 6.19 | 55.55 4.55 | |||||||||
| Vikhr 7B IT 5.4 | 51.91 8.89 | 55.00 12.58 | 53.32 10.58 | 47.66 10.86 | 44.44 8.49 | 45.88 9.46 | 49.60 9.75 | |||||||||
| VikhrGemma 2B IT | 45.93 4.26 | 39.17 6.72 | 42.13 5.50 | 42.09 3.57 | 49.07 5.93 | 45.22 4.01 | 43.68 3.91 | |||||||||
| AC | Gemma2 2B IT | 58.73 3.32 | 63.70 4.19 | 61.08 3.43 | 52.17 4.36 | 46.93 5.35 | 49.35 4.71 | 55.22 3.84 | ||||||||
| Gemma2 9B IT | 56.84 1.85 | 61.48 5.97 | 58.85 2.17 | 49.07 2.00 | 44.30 8.78 | 46.21 5.61 | 52.53 2.19 | |||||||||
| Qwen2 7B IT | 57.36 3.77 | 57.04 5.97 | 57.15 4.76 | 49.73 4.74 | 50.00 4.02 | 49.81 4.08 | 53.48 4.25 | |||||||||
| SaigaLlama3 8B | 59.02 3.26 | 65.56 5.70 | 62.03 3.81 | 53.16 4.52 | 46.05 5.84 | 49.19 4.61 | 55.61 3.74 | |||||||||
| Vikhr 7B IT 0.4 | 59.62 3.49 | 60.37 8.36 | 59.78 5.60 | 52.72 4.82 | 51.75 6.56 | 52.00 4.61 | 55.89 4.23 | |||||||||
| Vikhr 7B IT 5.4 | 61.87 2.07 | 62.96 8.18 | 62.13 4.30 | 55.50 3.03 | 53.95 7.25 | 54.33 3.58 | 58.23 2.38 | |||||||||
| VikhrGemma 2B IT | 57.77 2.90 | 64.81 6.21 | 60.98 3.77 | 51.40 4.47 | 43.86 6.39 | 47.11 4.91 | 54.05 3.57 |